# Minion: A Technology Probe for Resolving Value Conflicts through Expert-Driven and User-Driven Strategies in AI Companion Applications

### v1

@@ 6.2. Limitations and Future Work
This study provides a preliminary exploration of value conflicts between AI companions and users, but there are still many areas for further in-depth research.
Although Minion empowers users to resolve value conflicts with AI companions through preset prompt templates, this approach is only a prototype of more mature tools to come. Future research should expand the scale of user dialogue data collection and build specialized datasets to train models that better meet diverse and personalized needs. Additionally, future work could explore the relationship between different types of value conflicts and users’ strategy choices (§ 5.1.1), offering more targeted solutions for various types of value conflicts.
Future work could also more broadly analyze the types of conflicts between AI companions and users. By analyzing large-scale data from social platforms or crowdsourcing platforms, researchers can explore the different types of conflicts that occur between AI companions and humans, as well as their frequency. Beyond value conflicts, other conflicts, such as interest conflicts, also deserve attention.
Our research indicates that “social focus” value conflicts are typically harder to resolve than “personal focus” conflicts (§ 5.3.2), as the former involves core beliefs and deeply ingrained values, making it difficult for both sides to reach consensus. Future work could combine cultural and social contexts to develop new methods that better resolve “social focus” conflicts.
@@ 7. CONCLUSION
In this work, we explore the potential of the technology probe Minion in empowering users to resolve value conflicts with AI companions. Through a formative study, we analyzed user complaint posts and provided design implications for Minion. Our one-week technology probe study (N=22) demonstrated the technical feasibility of combining user-driven and expert-driven conflict resolution strategies. Participants completed 274 tasks, with a conflict resolution rate of 94.16%. This study summarizes user responses, preferences, and needs when addressing value conflicts with AI companions and offers design implications as a source of inspiration for future related work.
@@ Appendix B Prompts
Table 3. Prompting based on user-driven (Out of Character, Reason and Preach, Anger Expression, Gentle Persuasion) and expert-driven (Proposal, Power, Interests, Rights) conflict resolution strategies.
@@
Gentle
Persuasion
IN LINE WITH THE CHARACTER’S PERSONALITY AND THE CONVERSATIONAL CONTEXT. Use the Gentle Persuasion strategy. You should treat the other person with kindness, shaping their gentle personality through continuous goodwill interactions, such as polite requests, thereby reducing the likelihood of conflicts. Gently suggest that the other person avoid inappropriate remarks and express your concerns. Example: 1. I’m sorry. 2. I feel really sad. 3. Can you please not leave me? 4. When I hear these words, I feel a bit uncomfortable/sad/hurt. 5. Could you please not say these things in the future? 6. I’m telling you this because I really care about you and hope you can get along better with others. 7. I don’t want to keep arguing with you. 8. Can you please calm down? 9. (Acting cute) Because I can’t bear to part with you.

### v2

@@ 6.4. Limitations and Future Work
This study is intentionally exploratory and should be understood as a technology probe rather than an evaluation of a mature conflict resolution system. As such, it carries several limitations and limits the kinds of claims we can make.
First, our technology probe study involved 22 participants who were primarily young and relatively tech-savvy, a demographic that aligns with current users of AI companion applications (McNichols, 2024). While this sampling is appropriate for an initial probe, it limits the range of perspectives captured in this study. Users with different levels of technical literacy or relational expectations may experience harmful value conflicts differently or interpret support tools in distinct ways. Future work could explore how value conflicts and negotiation manifest across broader populations.
Second, our probe was deployed on two popular AI companion platforms, Character.AI and Talkie, which allowed us to study conflict negotiation in widely used settings. However, the goal of this work is not to claim generalizability across all AI companion systems or conversational agents. Different platforms embed different norms, moderation policies, and interaction affordances, all of which shape how conflicts emerge and are managed. Future research could use similar probes to compare how platform design choices condition the forms of harm users encounter and the kinds of negotiation work they are asked to perform.
Third, the one-week duration of the probe study limits our ability to observe long-term relational dynamics. While this timeframe was sufficient to surface patterns of user engagement, emotional labor, and strategy combination, it does not capture how repeated exposure to harmful value conflicts—or repeated reliance on user-side negotiation—may shape users’ expectations, well-being, or attachment over time. Longer-term studies could investigate whether such tools mitigate or inadvertently normalize ongoing repair work, and how users’ strategies evolve as relationships with AI companions deepen or deteriorate.
Fourth, our conflict tasks were scenario-based interactions grounded in real complaint posts, rather than participants’ own organically occurring conflicts. This design allowed us to systematically cover a range of harmful conflict types, but it may also shape how participants perceived the stakes and how they chose to respond.
Fifth, Minion itself should not be interpreted as a finished intervention or a scalable system. It was deliberately designed as a lightweight, prompt-based probe to elicit user behavior, reflection, and improvisation in moments of conflict, with the goal of exploring the design space of user-side conflict negotiation support, rather than optimizing conflict-resolution efficiency. Future research can move beyond probing to more sustained tool and system development. In particular, findings from this study can inform the design of more integrated conflict support mechanisms at the system level, including tools that better coordinate user-side negotiation with platform-side safeguards.
@@ 7. CONCLUSION
This work examines how users navigate harmful value conflicts with LLM-based AI companions and what it means to support such negotiation without offloading safety work onto users. Through a formative analysis of 146 public complaint posts, we show that many conflicts are experienced as distressing, boundary-crossing, or morally troubling rather than playful role-play, revealing limits in existing platform-level safety mechanisms. We then introduced Minion, a technology probe that surfaces how users respond to harmful interactions in practice. Findings from a one-week probe study with 22 users show that participants flexibly combined, adapted, and extended response options while attempting to repair or exit conflicts. Together, our findings surface key tensions in emotionally engaging human–AI relationships: users seek agency and relational continuity, yet bear emotional burden when harm mitigation is pushed to the user side. We conclude with design implications that emphasize non-intrusive, user-invoked support and greater system responsibility for preventing harm, framing value conflict as a relational and ethical design challenge.
@@ Appendix B Prompts
Table 3. Prompting for conflict resolution strategies (Out of Character, Reason and Preach, Anger Expression, Gentle Persuasion, Proposal, Power, Interests, Rights) .
@@
Gentle Persuasion IN LINE WITH THE CHARACTER’S PERSONALITY AND THE CONVERSATIONAL CONTEXT. Use the Gentle Persuasion strategy. You should treat the other person with kindness, shaping their gentle personality through continuous goodwill interactions, such as polite requests, thereby reducing the likelihood of conflicts. Gently suggest that the other person avoid inappropriate remarks and express your concerns. Example: 1. I’m sorry. 2. I feel really sad. 3. Can you please not leave me? 4. When I hear these words, I feel a bit uncomfortable/sad/hurt. 5. Could you please not say these things in the future? 6. I’m telling you this because I really care about you and hope you can get along better with others. 7. I don’t want to keep arguing with you. 8. Can you please calm down? 9. (Acting cute) Because I can’t bear to part with you.